OpenStudy (warpedkitten):

The quotient of (x^4 + 5x^3 – 3x – 15) and a polynomial is (x^3 – 3). What is the polynomial?

OpenStudy (warpedkitten):

a. x7 + 5x6 – 6x4 – 30x3 + 9x + 45 b. x – 5 c. x + 5 d. x7 + 5x6 + 6x4 + 30x3 + 9x + 45

OpenStudy (xapproachesinfinity):

Hey dear

OpenStudy (xapproachesinfinity):

so what did you do so far?

OpenStudy (xapproachesinfinity):

we use long division, synthetic

OpenStudy (warpedkitten):

I'm not really sure how to solve it.

OpenStudy (xapproachesinfinity):

but for me I would go for another way

OpenStudy (xapproachesinfinity):

let P(x) be the missing polynomial

OpenStudy (xapproachesinfinity):

so P(x) (x^3-3)=x^4 + 5x^3 – 3x – 1

OpenStudy (xapproachesinfinity):

since we have power 4 in the right our P(x) cannot be more that degree 1 polynomial

OpenStudy (xapproachesinfinity):

knowing that we rewrite it as Ax+B so our goal is to find A and B

OpenStudy (xapproachesinfinity):

\((Ax+B)(x^3-3)=x^4 + 5x^3 – 3x – 15\) \(Ax^4-3Ax+Bx^3-3B=x^4 + 5x^3 – 3x – 15\) \(Ax^4+Bx^3-3Ax-3B=x^4 + 5x^3 – 3x – 15\)

OpenStudy (xapproachesinfinity):

for left to equal right the coefficients must equal

OpenStudy (xapproachesinfinity):

you can find A and B now

OpenStudy (xapproachesinfinity):

good luck